OpenStudy (loveformusic1):

How do you find intervals of monotonicity and local extrema of functions? f(x)= x^3-x

OpenStudy (jskhupmang):

right find the derivative of f(x)

OpenStudy (jskhupmang):

then set it equal to zero

OpenStudy (jskhupmang):

and u will get extrema and test it whether the value is increasing or decreasing if f'(x)= 0 then the test fail

OpenStudy (jskhupmang):

aight u not even paying attention lol bye :D

OpenStudy (loveformusic1):

I'm sorry i'm new to this and didn't realize I had a response.

OpenStudy (mathmale):

"Local extrema" would include relative minima and relative maxima. (Forget about "absolute min. and max" for now. Please take the derivative of the given function and set the derivative = to 0. Show your work. What results do you obtain?

OpenStudy (loveformusic1):

How do I show you my work?

OpenStudy (mathmale):

Must be interesting exploring OpenStudy, given that you're new. Need your attention, tho.' Type it in: for example, f(x)=x^3-x, or \[f(x)=x^3-x\] Or draw it:

OpenStudy (mathmale):

Or take a screen shot and up load the image to OpenStudy.

OpenStudy (loveformusic1):

Okay sorry still trying to figure this out. Working on the problem now.

OpenStudy (loveformusic1):

so it would be 3x^2-1 for the derivative?

OpenStudy (loveformusic1):

3x^2-1=0 3x^2=1 x^2=(1/3) x= square root of (1/3)?

OpenStudy (mathmale):

Yes, but there are actually TWO roots. One is \[+\sqrt{\frac{ 1 }{ 3 }}.\] The other is \[-\sqrt{\frac{ 1 }{ 3}}.\]

OpenStudy (mathmale):

Please find the associated y value for each, using the original function.

OpenStudy (loveformusic1):

Why both positive and negative?

OpenStudy (mathmale):

First, you could square both of those results. In both cases you'd get +1/3. Secondly, an nth order polynomial has n roots. Your f '(x)=x^2-1 is a 2nd order poly. and thus has 2 roots. f(x)=x^3-x is a third order poly and has 3 roots. And so on.

OpenStudy (mathmale):

Given those 2 roots of f '(x) = 0, find the associated y-values, please.

OpenStudy (loveformusic1):

Oh okay you use the exponent? Okay one second please.

OpenStudy (mathmale):

The function is f(x)=x^3-x. Evaluate this at x=Sqrt(1/3). Your result? You might want to draw your result; it's a bit tricky to type it in.

OpenStudy (mathmale):

Here is what I'd type in myself:\[f(\frac{ 1 }{ \sqrt{3} })=(\frac{ 1 }{ \sqrt{3} })^3-\frac{ 1 }{ \sqrt{3} }\]

OpenStudy (mathmale):

This would be the value of the function at x=(1/Sqrt(3)).

OpenStudy (mathmale):

The two points you obtain this way are your local (or relative) extreme values.